Fringes found at channel 08 in hb-ke baseline, 08, 13 baseline hb-yg and none at ke-yg.

The watchdog timeouts on the log monitors no longer trigger an alarm. The message appears in the log monitor dialogue but doesn't do anything else (not even send an sms). (seen in both Hb and Yg). Addition, no alarm sets off a tone, even when “acknowledge alarm” becomes an option.


Module: HOB+0098, start pos: 3996.785 GB

iread power levels ifc and ifd are oscillating even though the attenuation isn't maxed. Jamie is investigating, although possibly just interference that can't be controlled.

Power dropping out a lot. Antenna needs re-connecting each time.

  • 17:00 UT - Clock drifted for second time, missed scans 291-1702 and 1705 to reconfig DBBC. (Ellen)
  • 19:50 UT - Clock drifted from about -2.4 to -3.2 us. Rebooted DBBC, no scans missed. (Ellen)
  • 12:15 UT - S RCP attenuator incorrectly set, presumably for whole experiment. HAS NOW BEEN SET TO 0dB. (Ross)


Module: HOB+0122, start pos 11405.548 GB

  • 17:00 UT - Experiment started ok. (Ellen)


Module: HOB+0126, start pos 9487.412 GB

Something happened to the fishlens camera between 8pm and 1am when I came back in - it's getting a very bad signal.

  • 17:00 UT - Experiment started ok. (Ellen)